# Decentralized Optimization with Coupled Constraints

### v3

@@ B.1 Dual problem
$\min_{x_{1},\ldots,x_{n}\in\mathbb{R}^{d}}\max_{w}\left[\sum_{i=1}^{n}f_{i}(x_{i})-\left\langle z,\mathbf{A}_{i}x_{i}\right\rangle\right]=\max_{w}\left[-\max_{x_{1},\ldots,x_{n}\in\mathbb{R}^{d}}\sum_{i=1}^{n}\left\langle\mathbf{A}_{i}^{\top}z,x_{i}\right\rangle-f_{i}(x_{i})\right]$
$\quad=-\min_{w}\sum_{i=1}^{n}f_{i}^{*}(\mathbf{A}_{i}^{\top}z).$
@@
$\min_{z_{1},\ldots,z_{n}}~{}$ $\sum_{i=1}^{n}g_{i}(z_{i}):=\sum_{i=1}^{n}f_{i}^{*}(\mathbf{A}_{i}^{\top}z_{i})$ (36)
[…]
B.2 Example graph
We follow the principle of lower bounds construction introduced in Kovalev et al. [2021] and take the example graph from Scaman et al. [2017]. Let the functions be held by the nodes be organized into a path graph with $n$ vertices, where $n$ is divisible by $3$. The nodes of graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ are divided into three groups $\mathcal{V}_{1}=\left\{1,\ldots,n/3\right\},\mathcal{V}_{2}=\left\{n/3+1,\ldots,2n/3\right\},\mathcal{V}_{3}=\left\{2n/3+1,\ldots,n\right\}$ of $n/3$ vertices each.
Now we recall the construction from Scaman et al. [2017]. Let $\gamma_{n}=\frac{1-\cos\left(\frac{\pi}{3n}\right)}{1+\cos\left(\frac{\pi}{3n}\right)}$. Since $\gamma_{n}\overset{n\to\infty}{\rightarrow}0$, there exists $n\geq 1$ such that $\gamma_{n}\geq\frac{1}{\chi}>\gamma_{n+1}$. Introduce edge weights $w_{i,i+1}=1-a\mathbb{I}\left\{i=1\right\}$, take the corresponding weighed Laplacian $W_{a}$ and denote its condition number $\chi$. If $a=1$, the network is disconnected and therefore $\chi(W_{a})=\infty$. If $a=0$, we have $\chi(W_{a})=1/\gamma_{n}$. By continuity of Laplacian spectra we obtain that for some $a\in[0,1)$ it holds $\chi(W_{a})=\chi$. Note that
$\gamma_{n}=\frac{1-\cos\frac{\pi}{3n}}{1+\cos\frac{\pi}{3n}}\leq\frac{\pi^{2}}{9n^{2}}~{}~{}\Rightarrow~{}~{}\chi\geq\frac{9n^{2}}{\pi^{2}}\geq n^{2}.$ (38)
B.3 Example functions
@@
$f_{i}(p,t)=\frac{\mu_{f}}{2}\left\|p-\frac{\sqrt{\hat{L}_{\mathbf{A}}}}{2\mu_{f}}e_{1}\right\|^{2}+\frac{L_{f}}{2}\|t\|^{2}.$
Correspondingly,
$f_{i}^{*}(u,v)=\frac{1}{2\mu_{f}}\|u\|^{2}+\frac{1}{2L_{f}}\|v\|^{2}-\frac{\sqrt{\hat{L}_{\mathbf{A}}}}{\mu_{f}}u_{1}.$
@@
$g_{i}(z)=f_{i}^{*}(\mathbf{A}_{i}^{\top}z)$ =$=\begin{cases}\frac{1}{2\mu_{f}}\|\sqrt{\hat{L}_{\mathbf{A}}}\mathbf{E}_{1}z\|^{2}+\frac{1}{2L_{f}}\|\sqrt{\hat{\mu}_{\mathbf{A}}}z\|^{2}-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{1}\\ 0,&i\in\mathcal{V}_{2}\\ \frac{1}{2\mu_{f}}\|\sqrt{\hat{L}_{\mathbf{A}}}\mathbf{E}_{2}z\|^{2}+\frac{1}{2L_{f}}\|\sqrt{\hat{\mu}_{\mathbf{A}}}z\|^{2}-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{3}\end{cases}$
=$=\begin{cases}\frac{1}{2}z^{\top}\left(\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}\mathbf{M}_{1}+\frac{\hat{\mu}_{\mathbf{A}}}{L_{f}}\mathbf{I}\right)z-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{1}\\ 0,&i\in\mathcal{V}_{2}\\ \frac{1}{2}z^{\top}\left(\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}\mathbf{M}_{2}+\frac{\hat{\mu}_{\mathbf{A}}}{L_{f}}\mathbf{I}\right)z-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{3}\end{cases}$ (39)
@@
$\sum_{i=1}^{n}g_{i}(z)$ $=\frac{n}{3}\left[\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z^{\top}(\mathbf{M}_{1}+\mathbf{M}_{2})z+\frac{\hat{\mu}_{\mathbf{A}}}{L_{f}}z^{\top}z-\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}z_{1}\right]$
$=\frac{n}{3}\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}\left[\frac{1}{2}z^{\top}\mathbf{M}z-z_{1}+\frac{\hat{\mu}_{\mathbf{A}}\mu_{f}}{\hat{L}_{\mathbf{A}}L_{f}}z^{\top}z\right],$
@@
B.4 Deriving the lower bound
Let us introduce the local memory $\mathcal{Z}_{i}(k)$ which is updated as follows.
1. If the algorithm performs a local gradient computation, then
$\mathcal{Z}_{i}(k+1)=\mathrm{Span}\left\{x,\nabla f_{i}(x),\nabla f_{i}^{*}(x):~{}x\in\mathcal{Z}_{i}(k)\right\}.$
2. If the algorithm performs a communication rounds, then
$\mathcal{Z}_{i}(k+1)=\mathrm{Span}\left\{x:~{}x\in\mathcal{Z}_{j}(k),~{}(i,j)\in\mathcal{E}\right\}.$
@@ 2. If the algorithm performs a communication rounds, then
$\rho=\frac{\sqrt{\frac{2}{3}\frac{L_{A}L_{f}}{\mu_{A}\mu_{f}}+1}-1}{\sqrt{\frac{2}{3}\frac{L_{A}L_{f}}{\mu_{A}\mu_{f}}+1}+1}.$
@@
$h(z)=\frac{1}{2}z^{\top}\mathbf{M}z+\frac{3\mu}{L-\mu}\|z\|^{2}-z_{1}$
@@
$\rho=\frac{\sqrt{\frac{2L}{3\mu}+\frac{1}{3}}-1}{\sqrt{\frac{2L}{3\mu}+\frac{1}{3}}+1}$
Let us deduce the expression for $L,\mu$ in terms of $L_{A},L_{f},\mu_{A},\mu_{f}$. We enforce $h(z)=\sum_{i=1}^{n}g_{i}(z)$ and get
$\frac{3\mu}{L-\mu}=\frac{\mu_{A}\mu_{f}}{L_{A}L_{f}}\Rightarrow\frac{L}{\mu}=1+\frac{L_{A}L_{f}}{\mu_{A}\mu_{f}}.$
Therefore, for $\rho$ we obtain
$\rho=\frac{\sqrt{\frac{2}{3}\frac{L_{A}L_{f}}{\mu_{A}\mu_{f}}+1}-1}{\sqrt{\frac{2}{3}\frac{L_{A}L_{f}}{\mu_{A}\mu_{f}}+1}+1}.$
Let us first show the lower bound on the number of communications.
@@
Let $s_{i}(k)$ denote the maximum index of a nonzero component of vector held by $i$-th node at step $k$, i.e.
$s_{i}(k)=\begin{cases}0,&\mathcal{Z}_{i}(k)\subseteq\left\{0\right\}\\ \min\left\{s\in\left\{1,2,\ldots\right\}:\mathcal{Z}_{i}(k)\subseteq\mathrm{Span}\left\{e_{1},\ldots,e_{s}\right\}\right\},&\text{else}.\end{cases}$
Let $k_{q}$ denote the the number of algorithm step by which exactly $q$ communication steps have been performed, where $q\geq 0$. For any $k\in\left\{1,\ldots,k_{q}\right\}$ we have
@@
Note that from the structure of $g_{i}(z)$, if the method performs a computation step, then
@@
Due to the structure of network, if the method makes a communication round, it follows
$s_{i}(k+1)\leq\begin{cases}\max\left(s_{i-1}(k),s_{i}(k),s_{i+1}(k)\right),&i\in\left\{2,\ldots,n-1\right\}\\ \max\left(s_{1}(k),s_{2}(k)\right),&i=1\\ \max\left(s_{n-1}(k),s_{n}(k)\right),&i=n\end{cases}$ (41)
@@ 2. For $q=(2\ell+1)(n/3+1),~{}\ell\in\left\{0,1,\ldots\right\}$ we have
Induction basis. Let $q=0$. From definitions of $g_{i}(z)$ it follows that
@@
Induction step for $q=(2\ell+1)(n/3+1)$. The proof follows by the same argument as for $q=(2\ell+1)(n/3+1)$.
@@
Now we estimate the distance to optimum.
$\|z_{i}(k)-z^{*}\|_{2}^{2}$ $\geq\sum_{\ell=s_{i}(k)+1}^{\infty}(z_{i}(k)-z^{*})^{2}=\sum_{\ell=s_{i}(k)+1}^{\infty}\rho^{2\ell}=\frac{\rho^{2s_{i}(k)+2}}{1-\rho^{2}}=\frac{\rho^{6+2\left\lfloor\frac{q}{n/3+1}\right\rfloor}}{1-\rho^{2}}$
$\overset{(a)}{\geq}\frac{\rho^{6+\frac{2q}{2n/3}}}{1-\rho^{2}}=\frac{\rho^{6}}{1-\rho^{2}}\cdot\rho^{\frac{3q}{n}}\overset{(b)}{=}\frac{\rho^{6}}{1-\rho^{2}}\cdot\rho^{\frac{3q}{\sqrt{\chi}}},$
@@
$\rho\geq\max\left(0,1-\sqrt{\frac{6\mu_{A}\mu_{f}}{L_{A}L_{f}}}\right).$
@@
$\|z_{i}(k)-z^{*}\|_{2}^{2}\geq\frac{\rho^{6}}{1-\rho^{2}}\left(\max\left(0,1-\sqrt{\frac{6\mu_{A}\mu_{f}}{L_{A}L_{f}}}\right)\right)^{\frac{3q}{\sqrt{\chi}}}.$
@@
$N_{\mathbf{W}}\geq\Omega\left(\sqrt{\chi}\sqrt{\frac{L_{A}L_{f}}{\mu_{A}\mu_{f}}}\log\left(\frac{1}{\varepsilon}\right)\right)$
and the number of oracle calls of $g_{i}(z)$ at each node (that is, the number of local matrix multiplications by $A_{i}$) is lower bounded as
$N_{\mathbf{A}}\geq\Omega\left(\sqrt{\frac{L_{A}L_{f}}{\mu_{A}\mu_{f}}}\log\left(\frac{1}{\varepsilon}\right)\right).$
B.5 Lower bound on the number of gradient computations

### v4

@@ C.1 Dual problem
$\min_{x_{1},\ldots,x_{n}\in\ell_{2}}\max_{z}\left[\sum_{i=1}^{n}f_{i}(x_{i})-\left\langle z,\mathbf{A}_{i}x_{i}\right\rangle\right]=\max_{z}\left[-\max_{x_{1},\ldots,x_{n}\in\ell_{2}}\sum_{i=1}^{n}\left\langle\mathbf{A}_{i}^{\top}z,x_{i}\right\rangle-f_{i}(x_{i})\right]$ (36)
$\quad=-\min_{z}\sum_{i=1}^{n}f_{i}^{*}(\mathbf{A}_{i}^{\top}z).$
@@
$\begin{aligned} \min_{z_{1},\ldots,z_{n}}~&\sum_{i=1}^{n}g_{i}(z_{i}):=\sum_{i=1}^{n}f_{i}^{*}(\mathbf{A}_{i}^{\top}z_{i})\\ &\text{s.t. }\mathbf{W}z=0.\end{aligned}$ (37)
C.2 Example graph
We follow the principle of lower bounds construction introduced in Kovalev et al. [2021] and take the example graph from Scaman et al. [2017]. Let the functions held by the nodes be organized into a path graph with $n$ vertices, where $n$ is divisible by $3$. The nodes of graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ are divided into three groups $\mathcal{V}_{1}=\left\{1,\ldots,n/3\right\},\mathcal{V}_{2}=\left\{n/3+1,\ldots,2n/3\right\},\mathcal{V}_{3}=\left\{2n/3+1,\ldots,n\right\}$ of $n/3$ vertices each.
Now we recall the construction from Scaman et al. [2017]. Maximum and minimum eigenvalues of a path graph have form $\lambda_{\max}(W)=2\left(1+\cos\frac{\pi}{n}\right),~\lambda_{\min^{+}}(W)=2\left(1-\cos\frac{\pi}{n}\right)$. Let $\beta_{n}=\frac{1+\cos\left(\frac{\pi}{n}\right)}{1-\cos\left(\frac{\pi}{n}\right)}$. Since $\beta_{n}\overset{n\to\infty}{\rightarrow}+\infty$, there exists $n=3m\geq 3$ such that $\beta_{n}\leq\kappa_{\mathbf{W}}<\beta_{n+3}$. For this $n$, introduce edge weights $w_{i,i+1}=1-a\mathbb{I}\left\{i=1\right\}$, take the corresponding weighed Laplacian $W_{a}$ and denote its condition number $\kappa(W_{a})$. If $a=1$, the network is disconnected and therefore $\kappa(W_{a})=\infty$. If $a=0$, we have $\kappa(W_{a})=\beta_{n}$. By continuity of Laplacian spectra we obtain that for some $a\in[0,1)$ it holds $\kappa(W_{a})=\kappa_{\mathbf{W}}$. Note that $\pi/(n+3)\in[0,\pi/3]$ for $x\in[0,\pi/3]$ it holds $1-\cos x\geq x^{2}/4$. We have
$\kappa_{\mathbf{W}}\leq\beta_{n+3}=\frac{1+\cos\frac{\pi}{n+3}}{1-\cos\frac{\pi}{n+3}}\leq\frac{72(n+3)^{2}}{\pi^{2}}\leq\frac{288n^{2}}{\pi^{2}}\leq 32n^{2}~~~\Rightarrow~~~\sqrt{\kappa_{\mathbf{W}}}\leq 4\sqrt{2}n=O(n).$ (38)
C.3 Example functions
@@
$f_{i}(p,t)=\frac{\mu_{f}}{2}\left\|p+\frac{\sqrt{\hat{L}_{\mathbf{A}}}}{2\mu_{f}}e_{1}\right\|^{2}+\frac{L_{f}}{2}\|t\|^{2}.$ (39)
We immediately note that each $f_{i}$ is $L_{f}$-smooth and $\mu_{f}$-strongly convex. The first term has the form $\frac{\mu_{f}}{2}\|p+ce_{1}\|^{2}$, and its convex conjugate is $\left(\frac{\mu_{f}}{2}\|p+ce_{1}\|^{2}\right)^{*}=\max_{p}\left\{\langle u,p\rangle-\frac{\mu_{f}}{2}\|p+ce_{1}\|^{2}\right\}$. The gradient by $p$ is $u-\mu_{f}(p+ce_{1})=0$, thus $p=\frac{1}{\mu_{f}}u-ce_{1}$ and $\max_{p}\left\{\langle u,p\rangle-\frac{\mu_{f}}{2}\|p+ce_{1}\|^{2}\right\}=\frac{\|u\|^{2}}{2\mu_{f}}-cu_{1}$. Correspondingly,
$f_{i}^{*}(u,v)=\frac{1}{2\mu_{f}}\|u\|^{2}+\frac{1}{2L_{f}}\|v\|^{2}-\frac{\sqrt{\hat{L}_{\mathbf{A}}}}{2\mu_{f}}u_{1}.$
@@
$g_{i}(z)=f_{i}^{*}(\mathbf{A}_{i}^{\top}z)$ $=\begin{cases}\frac{1}{2\mu_{f}}\|\sqrt{\hat{L}_{\mathbf{A}}}\mathbf{E}_{1}z\|^{2}+\frac{1}{2L_{f}}\|\sqrt{\hat{\mu}_{\mathbf{A}}}z\|^{2}-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{1}\\ 0,&i\in\mathcal{V}_{2}\\ \frac{1}{2\mu_{f}}\|\sqrt{\hat{L}_{\mathbf{A}}}\mathbf{E}_{2}z\|^{2}+\frac{1}{2L_{f}}\|\sqrt{\hat{\mu}_{\mathbf{A}}}z\|^{2}-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{3}\end{cases}$
$=\begin{cases}\frac{1}{2}z^{\top}\left(\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}\mathbf{M}_{1}+\frac{\hat{\mu}_{\mathbf{A}}}{L_{f}}\mathbf{I}\right)z-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{1}\\ 0,&i\in\mathcal{V}_{2}\\ \frac{1}{2}z^{\top}\left(\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}\mathbf{M}_{2}+\frac{\hat{\mu}_{\mathbf{A}}}{L_{f}}\mathbf{I}\right)z-\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z_{1},&i\in\mathcal{V}_{3}\end{cases}$ (40)
@@
$\begin{aligned} \sum_{i=1}^{n}g_{i}(z)&=\frac{n}{3}\left[\frac{\hat{L}_{\mathbf{A}}}{2\mu_{f}}z^{\top}(\mathbf{M}_{1}+\mathbf{M}_{2})z+\frac{\hat{\mu}_{\mathbf{A}}}{L_{f}}z^{\top}z-\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}z_{1}\right]\\ &=\frac{n}{3}\frac{\hat{L}_{\mathbf{A}}}{\mu_{f}}\left[\frac{1}{2}z^{\top}\mathbf{M}z-z_{1}+\frac{\hat{\mu}_{\mathbf{A}}\mu_{f}}{\hat{L}_{\mathbf{A}}L_{f}}z^{\top}z\right],\end{aligned}$ (41)
@@
C.4 Deriving the lower bound
@@
$\rho=\frac{\sqrt{\frac{2L_{\mathbf{A}}L_{f}}{\mu_{\mathbf{A}}\mu_{f}}+1}-1}{\sqrt{\frac{2L_{\mathbf{A}}L_{f}}{\mu_{\mathbf{A}}\mu_{f}}+1}+1}.$
@@
$h(z)=\frac{1}{2}z^{\top}\mathbf{M}z+\frac{3\mu}{L-\mu}z^{\top}z-z_{1}$
@@
$\rho=\frac{\sqrt{\frac{2L}{3\mu}+\frac{1}{3}}-1}{\sqrt{\frac{2L}{3\mu}+\frac{1}{3}}+1}.$
Let us deduce the expression for $L/\mu$ in terms of $L_{\mathbf{A}},L_{f},\mu_{\mathbf{A}},\mu_{f}$. We enforce $h(z)=\sum_{i=1}^{n}g_{i}(z)$ and set
$\frac{3\mu}{L-\mu}=\frac{\mu_{\mathbf{A}}\mu_{f}}{L_{\mathbf{A}}L_{f}}\Rightarrow\frac{L}{\mu}=1+3\frac{L_{\mathbf{A}}L_{f}}{\mu_{\mathbf{A}}\mu_{f}}.$
Therefore, $h(z)$ attains its minimum at $z^{k}=\rho^{k}$, where
$\rho=\frac{\sqrt{\frac{2L_{\mathbf{A}}L_{f}}{\mu_{\mathbf{A}}\mu_{f}}+1}-1}{\sqrt{\frac{2L_{\mathbf{A}}L_{f}}{\mu_{\mathbf{A}}\mu_{f}}+1}+1}.$
Let us first show the lower bound on the number of communications. Without loss of generality we can assume that the initial point chosen by a first-order algorithm is $x^{0}=0$ (otherwise we can shift the variables accordingly), thus $\mathcal{M}_{i}(0)=\{0\},~\mathcal{H}_{i}(0)=\left\{0\right\}$.
@@
Let $s_{i}(k)$ denote the maximum index of a nonzero component of vector blocks $p,t$ held by $i$-th node at step $k$ in its local memory $\mathcal{M}_{i}(k)$ and vector $z$ held by $i$-th node in its local memory $\mathcal{H}_{i}(k)$, i.e.
$s_{i}(k)=\begin{cases}0,&\hskip-28.45274pt\text{if }\mathcal{M}_{i}(k)\subseteq\left\{0\right\}\text{ and }\mathcal{H}_{i}(k)\subseteq\left\{0\right\}\\ \min\Big\{s\in\left\{1,2,\ldots\right\}:\mathcal{H}_{i}(k)\subseteq\mathrm{Span}\left\{e_{1},\ldots,e_{s}\right\}\\ \qquad~\text{ and }\mathcal{M}_{i}(k)\subseteq\mathrm{Span}\left\{e_{1},\ldots,e_{s}\right\}\times\left\{e_{1},\ldots,e_{s}\right\}\Big\},&\text{else}.\end{cases}$
Let $k_{q}$ denote the number of algorithm step by which exactly $q$ communication steps have been performed, where $q\geq 0$. For any $k\in\left\{1,\ldots,k_{q}\right\}$ we have
@@
If the method performs a multiplication by $\mathbf{A}_{i}$, it transfers the information from $\mathcal{M}_{i}(k)$ to $\mathcal{H}_{i}(k)$. If multiplication by $\mathbf{A}_{i}^{\top}$ is performed, the information is transferred in the opposite direction, i.e. from $\mathcal{H}_{i}(k)$ to $\mathcal{M}_{i}(k)$. If the method performs a matrix multiplication step (either by $\mathbf{A}$ or $\mathbf{A}^{\top}$), then from the structure of $\mathbf{A}_{i}$ we obtain
@@ 2. For $q=(2\ell+1)(n/3+1),~\ell\in\left\{0,1,\ldots\right\}$ we have
Induction basis. Let $q=0$. From definitions of $f_{i}$ and $\mathbf{A}_{i}$ it follows that
@@
Induction step for $q=2\ell(n/3+1)$. The proof follows by the same argument as for $q=(2\ell+1)(n/3+1)$.
@@
Now we estimate the distance to optimum. Due to the strong duality, the solution of problem (35) can be obtained from the solution of its dual (LABEL:prob:lower_dual) as
$x_{i}^{*}(z^{*})=\begin{pmatrix}p_{i}\\ t_{i}\end{pmatrix}(\mathbf{A}_{i}^{\top}z^{*})=\begin{pmatrix}\frac{\sqrt{\hat{L}_{\mathbf{A}}}}{\mu_{f}}(E_{(i)}z^{*}-e_{1})\\ \frac{\sqrt{\hat{\mu}_{\mathbf{A}}}}{L_{f}}z^{*}\end{pmatrix}$.
Therefore, denoting $\alpha=\frac{\hat{\mu}_{\mathbf{A}}}{L_{f}^{2}}$,
$\|x_{i}^{k}-x_{i}^{*}\|^{2}$ $\geq\|t_{i}^{k}-t_{i}^{*}\|^{2}\geq\sum_{\ell=s_{i}(k)+1}^{\infty}(t_{i,\ell}^{*})^{2}=\alpha\sum_{\ell=s_{i}(k)+1}^{\infty}(z_{\ell}^{*})^{2}=\alpha\sum_{\ell=s_{i}(k)+1}^{\infty}\rho^{2\ell}$
$=\alpha\frac{\rho^{2s_{i}(k)+2}}{1-\rho^{2}}=\alpha\frac{\rho^{6+2\left\lfloor\frac{q}{n/3+1}\right\rfloor}}{1-\rho^{2}}\overset{(a)}{\geq}\alpha\frac{\rho^{6+\frac{2q}{2n/3}}}{1-\rho^{2}}=\alpha\frac{\rho^{6}}{1-\rho^{2}}\cdot\rho^{\frac{3q}{n}}\overset{(b)}{\geq}\alpha\frac{\rho^{6}}{1-\rho^{2}}\cdot\rho^{\frac{12q\sqrt{2}}{\sqrt{\kappa_{\mathbf{W}}}}},$
@@
$\rho\geq\max\left(0,1-3\sqrt{\frac{2\mu_{\mathbf{A}}\mu_{f}}{L_{\mathbf{A}}L_{f}}}\right).$
@@
$\|x_{i}^{k}-x_{i}^{*}\|_{2}^{2}\geq\alpha\frac{\rho^{6}}{1-\rho^{2}}\left(\max\left(0,1-3\sqrt{\frac{2\mu_{\mathbf{A}}\mu_{f}}{L_{\mathbf{A}}L_{f}}}\right)\right)^{\frac{3q}{\sqrt{\kappa_{\mathbf{W}}}}}.$
@@
$N_{\mathbf{W}}\geq\Omega\left(\sqrt{\kappa_{\mathbf{W}}}\sqrt{\frac{L_{\mathbf{A}}L_{f}}{\mu_{\mathbf{A}}\mu_{f}}}\log\left(\frac{1}{\varepsilon}\right)\right),$ (45)
and the number of matrix multiplications at each node is lower bounded as
$N_{\mathbf{A}}\geq\Omega\left(\sqrt{\frac{L_{\mathbf{A}}L_{f}}{\mu_{\mathbf{A}}\mu_{f}}}\log\left(\frac{1}{\varepsilon}\right)\right).$ (46)
C.5 Lower bound on the number of gradient computations